# Geometry Aware Meta-Learning Neural Network for Joint Phase and Precoder Optimization in RIS

### v1

@@ Abstract
In reconfigurable intelligent surface (RIS) aided systems, the joint optimization of the precoder matrix at the base station and the phase shifts of the RIS elements involves significant complexity. In this paper, we propose a complex-valued, geometry aware meta-learning neural network that maximizes the weighted sum rate in a multi-user multiple input single output system. By leveraging the complex circle geometry for phase shifts and spherical geometry for the precoder, the optimization occurs on Riemannian manifolds, leading to faster convergence. We use a complex-valued neural network for phase shifts and an Euler inspired update for the precoder network. Our approach outperforms existing neural network-based algorithms, offering higher weighted sum rates, lower power consumption, and significantly faster convergence. Specifically, it converges faster by nearly 100 epochs, with a 0.7 bps improvement in weighted sum rate and a 1.8 dBm power gain when compared with existing work.
Index Terms:
reconfigurable intelligent surfaces, geometry, complex-valued neural network.
@@ I Introduction
Reconfigurable Intelligent Surfaces (RIS) will play an important role in emerging wireless systems and have been studied extensively in the last few years [1]. A key area of research has been focussing on the fact that in order to maximize the sum rate of the users, the precoder at the BS and the RIS phase shifts have to be optimized[2, 3, 4, 5]. The phase shifts of the RIS elements have to be optimized with the unit modulus constraint, while the precoder has to satisfy the transmit power constraint. The joint optimization problem is non-convex and NP-hard. Early approaches used alternating optimization where one of the optimization variables is kept fixed while the other gets optimized[6, 7, 8]. Specifically, the phase optimization is done using traditional iterative algorithms like Gradient Descent (GD) and the precoder optimization is inspired from the weighted sum MSE minimization (WMMSE) algorithm[9]. However, these approaches take longer to converge and involve costly matrix inversions, in addition to being sub-optimal. Hence recently neural networks were explored as a viable option.
@@
Deep reinforcement learning (DRL) based algorithms use deep deterministic policy gradient (DDPG) in order to perform the joint optimization[10, 5, 3]. In [3], a two-stage DRL based channel-oblivious algorithm is proposed. In [11], a deep learning based algorithm is proposed for a system with discretized phase shifts and imperfect channel state information (CSI). Following this work, the authors of [12] proposed a meta-learning based joint optimization framework using a shallow neural network which improved upon the results of [11]. An additional constraint on the search space of the precoder is enforced in [4] obtaining state-of-the-art improved sum rates.
@@
Although the neural network approaches have their merits, they do not exploit the inherent geometry of the optimization problem, namely, the underlying complex circle manifold corresponding to the phase shift constraint and the sphere manifold corresponding to the precoder power constraint. There are some non neural network based conventional methods which exploit complex circle manifold[13, 14, 15], however, they do not exploit the sphere manifold. A recent work (not neural network based) exploits both manifolds[16]. However, the corresponding precoder projection is highly computation intensive.
@@
In this paper, we design a complex-valued geometry aware meta-learning neural network where, (a) the weights are updated according to the Riemannian ADAM manifold optimizer, (b) both the complex circle geometry and the spherical geometry of the RIS phase shifts and the precoder entries respectively are leveraged, (c) we propose a complex-valued neural network since the complex circle manifold incorporation requires a complex space and (d) a recent Euler equation based update proposed in machine learning literature[17] is exploited for the precoder network design. Using the above complex-valued geometry aware neural network, we show significantly faster convergence and higher weighted sum rate against the state-of-the-art algorithm[4] for a multi-user multiple input single output (MU-MISO) system aided by an RIS. We name our approach as geometry aware meta learning neural network (GAMN).
@@ II System model
where $n_{k}\sim\mathcal{CN}(0,\sigma^{2})$ which is the additive complex Gaussian noise of variance $\sigma^{2}$. The first term of (1) denotes the signal corresponding to the target user while the second term denotes the interference term due to the other $K-1$ users. Hence the SINR at the $k^{th}$ user can be written as,
@@
The weighted sum rate is widely used as a performance metric in several works[9, 15, 11, 18, 14]. Using (2) it can be written as,
@@
P1: $\max\quad R(\mathbf{W},\boldsymbol{\Theta},\mathbf{H}_{BR},\mathbf{H}_{RU})$
@@ III-A Meta Learner
We then back-propagate the gradient of the loss and update the weights of the PL and the PRL. We denote the weights of the PL and the PRL as $\mathbf{x}_{P}$ and $\mathbf{x}_{PR}$ respectively. First, we take a closer look at P1 and the constraints of the problem. Let $\hat{\mathbf{w}}=[\mathbf{w}_{1}^{T},\ldots,\mathbf{w}_{K}^{T}]^{T}$. The constraint $\operatorname{tr}(\mathbf{W}^{H}\mathbf{W})\leq P$ can be written as $\hat{\mathbf{w}}^{H}\hat{\mathbf{w}}\leq P$. We define the manifold $\mathcal{M}_{\mathbf{W}}=\{\mathbf{x}\in\mathbb{R}^{2MK}:\mathbf{x}^{T}\mathbf{x}=1\}$ as a manifold embedded in $\mathbb{R}^{2MK}$. The tangent space can be found from the kernel of the differential $Dh(\mathbf{x})[\mathbf{v}]=\lim_{t\rightarrow 0}\frac{(\mathbf{x}+t\mathbf{v})^{T}(\mathbf{x}+t\mathbf{v})}{t}=0$ as $\mathbf{x}^{T}\mathbf{v}=0$. Hence we can write the tangent space $\mathcal{T}_{\mathbf{x}}\mathcal{M}_{W}$ as $\mathcal{T}_{\mathbf{x}}\mathcal{M}_{\mathbf{W}}=\{\mathbf{v}\in\mathbb{R}^{2MK}:\mathbf{v}^{T}\mathbf{x}=<\mathbf{x},\mathbf{v}>=0\}$. Now, $\mathcal{M}_{\mathbf{W}}$ is a Riemannian sub-manifold of $\mathbb{R}^{2MK}$. Hence the Riemannian gradient of a function $f:\mathcal{M}_{\mathbf{W}}\rightarrow\mathbb{R}$ at $\mathbf{x}\in\mathbb{R}^{2MK}$ can be found by, $\operatorname{grad}f(\mathbf{x})=\nabla f(\mathbf{x})-{\mathbf{x}}^{T}\nabla f(\mathbf{x})\mathbf{x}$. The retraction $\mathcal{R}_{\mathbf{W}}:\mathcal{T}_{\mathbf{x}}\mathcal{M}_{\mathbf{W}}\rightarrow\mathcal{M}_{\mathbf{W}}$ from the tangent space to the sphere manifold can be done by the operation
@@
Consider the manifold $\mathcal{M}_{\boldsymbol{\theta}}=\{x|x\in\mathbb{C},\left\lvert x\right\rvert=1\}$. It lies in a space defined by the equation $h(x)=\left\lvert x\right\rvert-1=0$. In order to find its tangent space we set the differential to zero as, $Dh(x)[v]=\lim_{t\rightarrow 0}\frac{(x+tv)^{*}(x+tv)-x^{*}x}{t}=0$ obtaining $v^{*}x=0$. Hence the tangent space is given by, $\mathcal{T}_{x}\mathcal{M}_{\boldsymbol{\theta}}=\{v|v^{*}x=0\}$. The Riemannian gradient of a function $f:\mathcal{M}_{\boldsymbol{\theta}}\rightarrow\mathbb{R}$ at can be found by, $\operatorname{grad}f(x)=\frac{\partial f(x)}{\partial x}-{x}^{*}\frac{\partial f(x)}{\partial x}x$. The retraction to the manifold $\mathcal{M}_{\boldsymbol{\theta}}$ is given by,
@@
Since the values of the phase shift matrix is constrained to be unit modulus and the corresponding values of the angles to lie in $[0,2\pi]$, we perform Riemannian optimization over $N$ complex circle manifolds corresponding to the $N$ RIS elements. Using the appropriate gradients, we perform Riemannian ADAM (RADAM) optimization according to the procedure in [19]. Hence now the optimization problem P1 becomes,
@@ III-B Phase Learner
We construct a complex-valued neural network consisting of a single hidden layer of $200$ neurons. The sizes of the input and the output layers are set equal to the number of RIS elements $N$. In [4], the authors show that the search for the local minima in the gradient space is more efficient compared to the search in the entire feasible region. Hence the input to the PL network consists of the gradient of the weighted sum rate with respect to the phase shifts $\boldsymbol{\theta}=[\theta_{1},\ldots,\theta_{N}]^{T}\in\mathbb{C}^{N\times 1}$ i.e, $\nabla_{\boldsymbol{\theta}}R(\mathbf{W},\boldsymbol{\Theta},\mathbf{H}_{BR},\mathbf{H}_{RU})$.
@@ III-C Precoder Learner
where we write $\mathbf{x}_{n}(t_{n})$ as $\mathbf{x}_{n}$. The Euler method was mapped to the success of robust Resnets in [17]. Hence we include the Euler factor in our precoder network. We perform the feed-forward operation for $n_{PR}$ times after which we do a projection to the sphere manifold according to (5). Note that in contrast to [16] where the retraction is done by the water-filling algorithm after every update, we ensure our iterates stay on the spherical manifold by exploiting Riemannian geometry. This helps our algorithm to converge significantly faster with improved sum rates.
We give a block diagram view of the Meta-learner (the outer loop minimizing the overall cost), the phase network and the precoder network in Fig. 1. We give our overall algorithm in Algorithm 1. While the precoder is updated in every iteration, the phase shifts are updated once in every $n_{I}$ iterations in order to maintain the stability similar to [4].
@@ IV Results
For our simulations, we set the same channel parameters and setup as in [4] and [12]. The channel between the BS and the RIS as well as between RIS and the users is assumed to follow Rician fading. Specifically,
@@
where $L_{RU_{k}}^{LoS}$, $L_{BR}^{LoS}$ and $L_{RU_{k}}^{NLoS}$, $L_{BR}^{NLoS}$ denote the path loss for the line of sight (LoS) and non line of sight (NLoS) components respectively. $\kappa_{RU}$ and $\kappa_{BR}$ are the Rician coefficients for the RIS-user and BS-RIS links respectively and are set equal to $10$. The BS is located at $(0,10)$ m, the users are randomly located in a circle of radius $5$ m centered at $(100,15)$ m, and the RIS at $(100,0)$ m. The path loss is set according to the 3GPP standard for LoS and NLoS components. The BS antenna spacing is set at $0.5\lambda$ with $28GHz$ frequency. We set the parameters $n_{M}$, $n_{P}$, $n_{PR}$, $\alpha_{P}$ and $\alpha_{PR}$ as $500$, $1$, $1$, $10^{-2}$ and $3.5\times 10^{-2}$ respectively. We also run the algorithm for $100$ channel realizations and average our results.
Figure 2: Weighted sum rate Vs. Power in dBm
We plot the weighted sum rate against the power in dBm in Fig. 2 for $N=100$. We compare the proposed GAMN algorithm against the state-of-the-art Gradient based manifold meta learning (GMML) algorithm proposed in [4]. In order to infer the impact of the complex-valued PL network, we plot results with the PL network having real valued weights which we denote as GAMNreal. Further, we include the Euler factor in all the algorithms mentioned above and tune them accordingly for the best results. From the plot, we infer that GAMN with the Euler factor set to $10$ gives the best results across the total power. Also, while the value of $h$ significantly improves the results of the GAMN algorithm, it does not improve GMML (i.e., we incorporated the Euler factor into GMML) or GAMNreal much. The proposed complex valued PL network when combined with the appropriately tuned Euler factor yields the best results, both across power and across the number of RIS elements $N$ which is shown in Fig. 3. We set the total power to be $10$ dBm for Fig. 3. We observe $1.8$ dBm of power gain from Fig. 2 comparing GAMN with $h=10$ and GMML. GAMN outperforms GMML by a weighted sum rate difference of $0.7$ bps with $10$ dBm power and $N=100$.
Figure 3: Weighted sum rate Vs. number of RIS elements $N$
We give the convergence plot of our algorithm and the GMML algorithm with the total power set at $10$ dBm and appropriately tuned Euler factor in Fig. 4. We observe that GAMN is able to converge faster than the GMML algorithm by an order of nearly $100$ epochs, owing to the manifold based optimization and the Euler factor we incorporate into our algorithm. Optimization over manifolds reduces the search space and speeds up convergence to the global minimum, which accounts for the faster convergence rate.

### v2

@@ Abstract
In reconfigurable intelligent surface (RIS) aided systems, the joint optimization of the precoder matrix at the base station and the phase shifts of the RIS elements involves significant complexity. In this paper, we propose a complex-valued, geometry aware meta-learning neural network that maximizes the weighted sum rate in a multi-user multiple input single output system. By leveraging the complex circle geometry for phase shifts and spherical geometry for the precoder, the optimization occurs on Riemannian manifolds, leading to faster convergence. We use a complex-valued neural network for phase shifts and an Euler inspired update for the precoder network. Our approach outperforms existing neural network-based algorithms, offering higher weighted sum rates, lower power consumption, and significantly faster convergence. Specifically, it converges faster by nearly 100 epochs, with a 0.7 bps improvement in weighted sum rate and a 1.8 dB power gain when compared with existing work. Further it outperforms the state-of-the-art alternating optimization algorithm by 0.86 bps with a 2.6 dB power gain.
@@ I Introduction
Reconfigurable Intelligent Surfaces (RIS) will play an important role in emerging wireless systems and have been studied extensively in the last few years [pan2021reconfigurable]. A key area of research has been focusing on the fact that in order to maximize the sum rate of the users, the precoder at the BS and the RIS phase shifts have to be optimized[subhash2023optimal, nayak2024drl, zhu2024robust, chen2023integrated]. The phase shifts of the RIS elements have to be optimized with the unit modulus constraint, while the precoder has to satisfy the transmit power constraint. The joint optimization problem is non-convex and NP-hard. Early approaches used alternating optimization where one of the optimization variables is kept fixed while the other gets optimized[huang2019reconfigurable, zhang2020sum, abeywickrama2020intelligent]. Specifically, the phase optimization is done using traditional iterative algorithms like Gradient Descent (GD) and the precoder optimization is inspired from the weighted sum MSE minimization (WMMSE) algorithm[shi2011iteratively]. However, these approaches take longer to converge and involve costly matrix inversions, in addition to being sub-optimal. Hence recently neural networks were explored as a viable option.
@@
Deep reinforcement learning (DRL) based algorithms use deep deterministic policy gradient (DDPG) in order to perform the joint optimization[huang2021multi, chen2023integrated, nayak2024drl]. In [nayak2024drl], a two-stage DRL based channel-oblivious algorithm is proposed. In [xu2021robust], a deep learning based algorithm is proposed for a system with discretized phase shifts and imperfect channel state information (CSI). Following this work, the authors of [wang2024] proposed a meta-learning based joint optimization framework using a shallow neural network which improved upon the results of [xu2021robust]. An additional constraint on the search space of the precoder is enforced in [zhu2024robust] obtaining state-of-the-art improved sum rates.
@@
Although the neural network approaches have their merits, they do not exploit the inherent geometry of the optimization problem, namely, the underlying complex circle manifold corresponding to the phase shift constraint and the sphere manifold corresponding to the precoder power constraint. There are some non neural network based conventional methods which exploit complex circle manifold[yu2019miso, li2020weighted, guo2020weighted], however, they do not exploit the sphere manifold. A recent work (not neural network based) exploits both manifolds[perovic2021achievable]. However, the corresponding precoder projection is highly computation intensive.
@@
In this paper, we design a complex-valued geometry aware meta-learning neural network where, (a) the weights are updated according to the Riemannian ADAM manifold optimizer, (b) both the complex circle geometry and the spherical geometry of the RIS phase shifts and the precoder entries respectively are leveraged, (c) we propose a complex-valued neural network since the complex circle manifold incorporation requires a complex space and (d) a recent Euler equation based update proposed in machine learning literature[zhang2019towards] is exploited for the precoder network design. Using the above complex-valued geometry aware neural network, we show significantly faster convergence and higher weighted sum rate against the state-of-the-art neural-network based algorithm[zhu2024robust] as well a state-of-the-art traditional alternating optimization (AO) approach [li2020weighted] for a multi-user multiple input single output (MU-MISO) system aided by an RIS. We name our approach as geometry aware meta learning neural network (GAMN).
@@ II System model
where $n_{k}\sim\mathcal{CN}(0,\sigma^{2})$ which is the additive complex Gaussian noise of variance $\sigma^{2}$. The first term of (1) denotes the signal corresponding to the target user while the second term denotes the interference term due to the other K1K-1 users. Hence the SINR at the $k^{th}$ user can be written as,
@@
The weighted sum rate is widely used as a performance metric in several works[shi2011iteratively, guo2020weighted, xu2021robust, xia2021meta, li2020weighted]. Using (2) it can be written as,
@@
P1: $\max_{\mathbf{W},\boldsymbol{\Theta}}\quad R(\mathbf{W},\boldsymbol{\Theta},\mathbf{H}_{BR},\mathbf{H}_{RU})$
@@ III-A Meta Learner
We then back-propagate the gradient of the loss and update the weights of the PL and the PRL. We denote the weights of the PL and the PRL as $\mathbf{x}_{P}$ and $\mathbf{x}_{PR}$ respectively. First, we take a closer look at P1 and the constraints of the problem. Let $\hat{\mathbf{w}}=[\mathbf{w}_{1}^{T},\ldots,\mathbf{w}_{K}^{T}]^{T}$. The constraint $\operatorname{tr}(\mathbf{W}^{H}\mathbf{W})\leq P$ can be written as $\hat{\mathbf{w}}^{H}\hat{\mathbf{w}}\leq P$. We define the manifold,
$\mathcal{M}_{\mathbf{W}}=\{\mathbf{x}\in\mathbb{R}^{2MK}:\mathbf{x}^{T}\mathbf{x}=1\},$
which is a manifold embedded in $\mathbb{R}^{2MK}$. The tangent space can be found from the kernel of the differential $Dh(\mathbf{x})[\mathbf{v}]=\lim_{t\rightarrow 0}\frac{(\mathbf{x}+t\mathbf{v})^{T}(\mathbf{x}+t\mathbf{v})}{t}=0$ as $\mathbf{x}^{T}\mathbf{v}=0$. Hence we can write the tangent space $\mathcal{T}_{\mathbf{x}}\mathcal{M}_{W}$ as,
$\mathcal{T}_{\mathbf{x}}\mathcal{M}_{\mathbf{W}}=\{\mathbf{v}\in\mathbb{R}^{2MK}:\mathbf{v}^{T}\mathbf{x}=<\mathbf{x},\mathbf{v}>=0\}.$
Now, $\mathcal{M}_{\mathbf{W}}$ is a Riemannian sub-manifold of $\mathbb{R}^{2MK}$. Hence the Riemannian gradient of a function $f:\mathcal{M}_{\mathbf{W}}\rightarrow\mathbb{R}$ at $\mathbf{x}\in\mathbb{R}^{2MK}$ can be found by, $\operatorname{grad}f(\mathbf{x})=\nabla f(\mathbf{x})-{\mathbf{x}}^{T}\nabla f(\mathbf{x})\mathbf{x}$. The retraction $\mathcal{R}_{\mathbf{W}}:\mathcal{T}_{\mathbf{x}}\mathcal{M}_{\mathbf{W}}\rightarrow\mathcal{M}_{\mathbf{W}}$ from the tangent space to the sphere manifold can be done by the operation
@@
Consider the complex circle manifold given by,
$\mathcal{M}_{\boldsymbol{\theta}}=\{x|x\in\mathbb{C},\left\lvert x\right\rvert=1\}.$
It lies in a space defined by the equation $h(x)=\left\lvert x\right\rvert-1=0$. In order to find its tangent space we set the differential to zero as, $Dh(x)[v]=\lim_{t\rightarrow 0}\frac{(x+tv)^{*}(x+tv)-x^{*}x}{t}=0$ obtaining $v^{*}x=0$. Hence the tangent space is given by,
$\mathcal{T}_{x}\mathcal{M}_{\boldsymbol{\theta}}=\{v|v^{*}x=0\}.$
The Riemannian gradient of a function $f:\mathcal{M}_{\boldsymbol{\theta}}\rightarrow\mathbb{R}$ can be found by, $\operatorname{grad}f(x)=\frac{\partial f(x)}{\partial x}-{x}^{*}\frac{\partial f(x)}{\partial x}x$. The retraction to the manifold $\mathcal{M}_{\boldsymbol{\theta}}$ is given by,
@@
Since the values of the phase shift matrix are constrained to be unit modulus and the corresponding values of the angles to lie in $0,2\pi]$, we perform Riemannian optimization over $N$ complex circle manifolds corresponding to the $N$ RIS elements. Using the appropriate gradients, we perform Riemannian ADAM (RADAM) optimization according to the procedure in [becigneulriemannian]. Hence now the optimization problem P1 becomes,
@@ III-B Phase Learner
We construct a complex-valued neural network consisting of a single hidden layer of $200$ neurons. The sizes of the input and the output layers are set equal to the number of RIS elements $N$. In [zhu2024robust], the authors show that the search for the local minima in the gradient space is more efficient compared to the search in the entire feasible region. Hence the input to the PL network consists of the gradient of the weighted sum rate with respect to the phase shifts $\boldsymbol{\theta}=[\theta_{1},\ldots,\theta_{N}]^{T}\in\mathbb{C}^{N\times 1}$ i.e, $\nabla_{\boldsymbol{\theta}}R(\mathbf{W},\boldsymbol{\Theta},\mathbf{H}_{BR},\mathbf{H}_{RU})$.
@@ III-C Precoder Learner
where we write $\mathbf{x}_{n}(t_{n})$ as $\mathbf{x}_{n}$. The Euler method was mapped to the success of robust Resnets in [zhang2019towards]. Hence we include the Euler factor in our precoder network. We perform the feed-forward operation for $n_{PR}$ times after which we do a projection to the sphere manifold according to (5). Note that in contrast to [perovic2021achievable] where the retraction is done by the water-filling algorithm after every update, we ensure our iterates stay on the spherical manifold by exploiting Riemannian geometry. This helps our algorithm to converge significantly faster with improved sum rates.
We give a block diagram view of the Meta-learner (the outer loop minimizing the overall cost), the phase network and the precoder network in Fig. 1. We give our overall algorithm in Algorithm 1. While the precoder is updated in every iteration, the phase shifts are updated once in every $n_{I}$ iterations in order to maintain the stability similar to [zhu2024robust].
@@ IV Results
For our simulations, we set the same channel parameters and setup as in [zhu2024robust] and [wang2024]. The channel between the BS and the RIS as well as between RIS and the users is assumed to follow Rician fading. Specifically,
@@
where $L_{RU_{k}}^{LoS}$, $L_{BR}^{LoS}$ and $L_{RU_{k}}^{NLoS}$, $L_{BR}^{NLoS}$ denote the path loss for the line of sight (LoS) and non line of sight (NLoS) components respectively. $\kappa_{RU}$ and $\kappa_{BR}$ are the Rician coefficients for the RIS-user and BS-RIS links respectively and are set equal to $10$. The BS is located at $(0,10)$ m, the users are randomly located in a circle of radius $5$ m centered at $(100,15)$ m, and the RIS at $(100,0)$ m. The path loss is set according to the 3GPP standard for LoS and NLoS components [3gpp]. The BS antenna spacing is set at $0.5\lambda$ with $28GHz$ frequency. We set the parameters $n_{M}$, $n_{P}$, $n_{PR}$, $\alpha_{P}$ and $\alpha_{PR}$ as $500$, 11, 11, $10^{-2}$ and $3.5\times 10^{-2}$ respectively. We also run the algorithm for $100$ channel realizations and average our results. The default values of power, N, and M, unless otherwise specified, are 10 dB, 100, and 64, respectively.
@@
In Fig. 2, we show the convergence behavior of our algorithm and compare it with the GMML algorithm, with the total power set to 10 dBm and the Euler factor appropriately tuned. We observe that GAMN converges nearly 100 epochs faster than GMML, owing to the manifold-based optimization and the incorporation of the Euler factor in our algorithm. Optimization on manifolds reduces the search space and accelerates convergence toward the global minimum, which explains the improved convergence rate. The traditional optimization technique for this problem employs an AO approach, where the beamforming matrix is optimized using the WMMSE algorithm and the phase shifts are optimized via the Riemannian conjugate gradient (RCG) algorithm [li2020weighted]. This results in a multi-stage iterative process in which each subproblem requires its own iterative update procedure. In particular, the WMMSE algorithm involves matrix inversions in every update step, further increasing the computational complexity. The per-iteration computational cost and therefore the overall time complexity of AO is not comparable to that of GAMN or GAMML. Consequently, we do not include a convergence plot for AO in this setting. Note that AO converges within 100 iterations to a value of 10.0844, whereas GMML and GAMN surpass this value at epochs 200 and 100, respectively, with final converged values of 10.2594 and 10.9526.
Figure 3: Weighted Sum Rate v/s Power
In Fig. 3, we plot the weighted sum rate versus the total transmit power for $N=100$ and $M=64$. We compare the proposed GAMN algorithm with the state-of-the-art Gradient-based Manifold Meta-Learning (GMML) algorithm proposed in [zhu2024robust], as well as the traditional alternating optimization method [li2020weighted]. To evaluate the impact of the complex-valued PL network, we also include a version of GAMN in which the PL network uses real-valued weights, denoted as GAMNreal. Furthermore, we incorporate the Euler factor into all the aforementioned algorithms and tune their parameters for optimal performance. From the plot, we observe that GAMN with the Euler factor set to $10$ achieves the best performance across all power levels. In addition, although the value of $h$ significantly improves the performance of GAMN, it does not yield substantial improvements for GMML or GAMNreal.
Figure 4: Weighted Sum Rate v/s number of RIS elements
Figure 5: Weighted Sum Rate v/s number of BS antennas
In Fig. 4 and Fig. 5, we plot the weighted sum rate versus the number of RIS elements ($N$) and the number of BS antennas ($M$). We clearly observe that the proposed GAMN method with an Euler factor of 10 outperforms both the state-of-the-art GMML method and the traditional optimization approach. Moreover, the proposed complex-valued PL network, when combined with an appropriately tuned Euler factor, achieves the best performance for different values of the number of RIS elements $N$ and the number of BS antennas $M$.
We observe $1.8$ dB of power gain from Fig. 3 comparing GAMN with $h=10$ and GMML. GAMN outperforms GMML by a weighted sum rate difference of $0.7$ bps with $10$ dBm power. Furthermore, GAMN outperforms the traditional AO baseline by $0.86$ bps at $10$ dBm power and a power gain of $2.6$ dB.
Complexity analysis: The computational cost of the SINR evaluation is dominated by the term $h_{k}^{H}\Theta Gw_{j}$, whose computation requires $O(K^{2}MN)$ multiplications. When this cost is combined with the iterative structure of the algorithm, where $N_{e}$ denotes the number of epochs, $N_{o}$ the number of outer iterations per epoch, and $N_{i}$ the number of inner iterations within each outer iteration, the total computational complexity becomes $O(N_{e}N_{o}N_{i}K^{2}MN)$, which is the same as for [zhu2024robust]. Our choice of optimizer introduces no further computational overhead beyond this dominant term. Importantly, no cubic-order terms appear in the complexity expression, since the method does not rely on matrix inversions as in alternating optimization [li2020weighted].